# cosmICweb: Cosmological Initial Conditions for Zoom-in Simulations in the Cloud

### v1

@@ V.3 Example Simulations
Figure 11: Comparison of the dark matter density and gas properties between three re-simulations of the halo described in Richings et al. (2021) using swift (Schaller et al., 2024) (top), gizmo (Hopkins, 2015) (middle), and ramses (Teyssier, 2002) (bottom). The four columns show the dark matter density, gas density, gas temperature, and gas entropy in a slice perpendicular to the $z$-axis centered on the halo. The white circle indicates the $R_{200c}$ spherical overdensity radius measured in the gizmo run. See main text for simulation details.
@@
The second halo we resimulate is a group-mass halo that has been featured in a study quantifying the impact of baryonic physics on halo and substructure properties (Richings et al., 2021). This halo is also from the Eagle 100 Mpc volume and its GroupID from the Eagle database in the reference model is 129 in snapshot 26 (z=0.18z=0.18) [^8] Initial conditions on cosmICweb can be accessed under https://cosmicweb.eu/publication/group-sized-halo-richings-2021.
The initial conditions for the resimulation have been generated with music using levelmin=7, levelmin_TF=10, and levelmax=11, with the zoom region obtained from the $R_{tb}=8R{200c}$ Lagrangian ellipsoid. For these resimulations, we follow a different strategy in that we demonstrate how cosmICweb facilitates the reproducibility of cosmological simulation results across different simulation codes by leveraging the plethora of output plugins available for the music code. Specifically, we show resimulations, at fixed resolution of the initial conditions, using the ramses (Teyssier, 2002) code (Eulerian adaptive mesh finite volume hydrodynamics and adaptive mesh particle mesh gravity solver), the gizmo (Hopkins, 2015) code (Mesh-free Lagrangian finite volume hydrodynamics with a tree-particle mesh gravity solver), and swift (Schaller et al., 2024) (smooth particle hydrodynamics (SPH) Lagrangian hydrodynamics with fast multipole-particle mesh gravity solver) code. For these simulations, initial conditions were generated for both dark matter and baryons (neglecting any baryon isocurvature perturbations so that the two initially perfectly trace each other). We assumed non-radiative evolution for the gas, disabling any extra astrophysics such as a UV background and feedback processes. For gizmo and swift, we set a comoving gravitational force softening of $2\,\text{kpc}$. For ramses, we enabled Lagrangian refinement with a maximum refinement level of 18, so that the smallest grid cells at $z=0$ have a linear size of $\sim 0.4$ kpc.
@@ VI Conclusion
We provide an updated version 2 of the music  [^9] The updated version 2 of music is available from https://github.com/cosmo-sims/MUSIC initial condition generator software that supports Panphasia noise fields along with other improvements such as e.g., a reduced memory usage and direct interfacing with the Class linear Einstein-Boltzmann code.
Development of cosmICweb is a continuing effort, since the data as well as the requirements change over time, requiring adaptations and extensions to the project. We categorize future developments into two classes: the addition of new data, in particular the integration of new simulations, and the addition of new features. The kind of improvements that we will consider highly depends on the feedback and requirements from the community.
VI.1 New simulations and data
New halo catalogs and proto-halos from additional simulations can either be directly hosted on cosmICweb servers, or from existing or new databases located outside. To facilitate the setting up of the API interface, we may publish our own implementation described in Section IV.3 as a template in the future.
VI.2 Additional functionality
@@ Acknowledgements
Appendix A API Server Requirements
@@
Appendix B cosmicweb-music Command Line Interface

### v3

@@ 5.2 Group-sized halo
Figure 11: Comparison of the dark matter density and gas properties between three re-simulations of the halo described in Richings et al. (2021) using swift (Schaller et al., 2024) (top), gizmo (Hopkins, 2015) (middle), and ramses (Teyssier, 2002) (bottom). The four columns show the dark matter density, gas density, gas temperature, and gas entropy in a slice perpendicular to the $z$-axis centered on the halo. The white circle indicates the $R_{200,c}$ spherical overdensity radius measured in the gizmo run. See main text for simulation details.
@@
The second halo we resimulate is a group-mass halo that has been featured in a study quantifying the impact of baryonic physics on halo and substructure properties (Richings et al., 2021). This halo is also from the Eagle 100 Mpc volume and its GroupID from the Eagle database in the reference model is 129 in snapshot 26 ($z=0.18$) [^9] Initial conditions on cosmICweb can be accessed under https://cosmicweb.eu/publication/group-sized-halo-richings-2021.
[…]
@@ 6 Conclusion
[…]
Development of cosmICweb is a continuing effort, since the data as well as the requirements change over time, requiring adaptations and extensions to the project. The kind of improvements that we will consider highly depends on the feedback and requirements from the community. We outline our plans in Appendix A.
@@ Acknowledgements
Appendix A Future Directions
We categorize future developments into two classes: the addition of new data, in particular the integration of new simulations, and the addition of new features.
A.1 New simulations and data
New halo catalogs and proto-halos from additional simulations can either be directly hosted on cosmICweb servers, or from existing or new databases located outside. To facilitate the setting up of the API interface, we may publish our own implementation described in Section 4.3 as a template in the future.
A.2 Additional functionality
Appendix B API Server Requirements
@@
Appendix C cosmicweb-music Command Line Interface